OpenStudy (anonymous):

what is 0.08 as a fraction and a percentage

OpenStudy (anonymous):

2 over 25 and 8%

OpenStudy (anonymous):

The percentage would be 8%

OpenStudy (igreen):

0.08 is the same thing as 8/100, but you can simplify that by dividing 4 to 8 and 100.

OpenStudy (igreen):

@Boing!AnImeFanGirl @DragonOrca97 Don't give answers. Read the Code of Conduct: http://openstudy.com/code-of-conduct

OpenStudy (skullpatrol):

$$\Huge 0.08=\frac{8}{100}=8\cdot\frac{1}{100}=8\%$$

OpenStudy (skullpatrol):

because 1/100 = %

OpenStudy (igreen):

@skullpatrol Don't give answers.

OpenStudy (anonymous):

To make that easier you just take 0.08 times 100 and then you get your answer!

OpenStudy (skullpatrol):

I am not giving the answer @iGreen by definition percent means "hundredths" or "divide by 100"

OpenStudy (skullpatrol):

"divide by 100" is the same as multiplying by 1/100 so rewriting the number as x * (1/100) gives x %

OpenStudy (igreen):

@skullpatrol You just gave her the answer '8%'.

OpenStudy (skullpatrol):

So did you @iGreen when you wrote "0.08 is the same thing as 8/100"

OpenStudy (anonymous):

dividing by 100 is not the same. when you multiply by 100 you get the percentage. when you divide by 100 you get the decimal form. she was asking for the percentage so then that means she will use x times 100 to get her answer.

OpenStudy (skullpatrol):

0.08 is the same thing as 8/100 = 8 * (1/100) = 8 %

OpenStudy (skullpatrol):

there is no need to multiply by 100 @DragonOrca97 since 1/100 = %

OpenStudy (igreen):

8/100 isn't the answer -.-

OpenStudy (igreen):

You have to simplify it.

OpenStudy (anonymous):